OpenStudy (toxicsugar22):

can someone help me with this question log(x+3)+log (x)=1

OpenStudy (toxicsugar22):

sorry

OpenStudy (anonymous):

base 10?

OpenStudy (anonymous):

yes

OpenStudy (toxicsugar22):

log(x+3)+logx=1

OpenStudy (anonymous):

log[(x+3)*(x)]=1

OpenStudy (anonymous):

oh yeah. rewrite as one log \[log(x(x+3)) = 1\] so \[log(x^2+3x)=1\] then rewrite in equivalent experiential form \[x^2+3x=10^1\]

OpenStudy (anonymous):

now you have a quadratic to solve \[x^2+3x-10=0\] \[(x+5)(x-2)=0\] \[x+5=0\] \[x=-5\] or \[x-2=0\] \[x=2\]

OpenStudy (amistre64):

5 and 2 i think are better options ;)

OpenStudy (anonymous):

but you cannot take the log of a negative number so answer is 2

OpenStudy (anonymous):

no i like my answers. it does factor as \[(x+5)(x-2)\] right?

OpenStudy (anonymous):

yeah

OpenStudy (toxicsugar22):

ok thanks i got that

OpenStudy (anonymous):

whew so answer is 2

OpenStudy (anonymous):

don't put -5!

OpenStudy (anonymous):

yeah

OpenStudy (toxicsugar22):

i didnt

OpenStudy (amistre64):

5 and 2 are better options for the factoring than 5 and 3 :)

OpenStudy (toxicsugar22):

can u help me with this one to

OpenStudy (anonymous):

did i put 5 and 3?

OpenStudy (amistre64):

....no, the deleted post did tho

OpenStudy (anonymous):

no i did

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

you are paying too close attention!

OpenStudy (amistre64):

lol

OpenStudy (amistre64):

just trying to keep my mind of the money pit im driving in

OpenStudy (toxicsugar22):

so it is 2

OpenStudy (toxicsugar22):

the answer is 2

OpenStudy (toxicsugar22):

is it

OpenStudy (anonymous):

yes 2

OpenStudy (toxicsugar22):

and not -5

OpenStudy (anonymous):

because \[log(5)+log(2)=log(10)=1\]

OpenStudy (anonymous):

you cannot take the log of a negative number

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

if i write \[log(-5)=x\] that means \[10^x=-5\] but \[10^x\] is always positive so it cannot be -5

OpenStudy (toxicsugar22):

ok thanks

OpenStudy (anonymous):

welcome